### v1

# Disambiguate Entity Matching through Relation Discovery with Large Language Models

Abstract.
Entity matching is a critical challenge in data integration and cleaning, central to tasks like fuzzy joins and deduplication. Traditional approaches have focused on overcoming fuzzy term representations through methods such as edit distance, Jaccard similarity, and more recently, embeddings and deep neural networks, including advancements from large language models (LLMs) like GPT. However, the core challenge in entity matching extends beyond term fuzziness to the ambiguity in defining what constitutes a ”match,” especially when integrating with external databases. This ambiguity arises due to varying levels of detail and granularity among entities, complicating exact matches. We propose a novel approach that shifts focus from purely identifying semantic similarities to understanding and defining the ”relations” between entities as crucial for resolving ambiguities in matching. By predefining a set of relations relevant to the task at hand, our method allows analysts to navigate the spectrum of similarity more effectively, from exact matches to conceptually related entities.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Entity matching, also known as record linkage, is the fundamental task for performing fuzzy join for data integration (Chen et al., 2019) and deduplication for data cleaning (Manghi et al., 2020). Previous efforts concentrated on the fuzzy term representations, such as synonyms and abbreviations. To match fuzzy terms, traditional methods like edit distance and Jaccard similarity have been used to measure term similarity. To capture semantic similarities between terms, techniques such as embeddings (Lu et al., 2021) and deep neural networks (Barlaug and Gulla, 2021; Mudgal et al., 2018) have been utilized. Recently, large language models (LLMs) like GPT have achieved results that are comparable, or even better than, previous SOTA approaches, with zero or few shot learning prompts (Nuntachit and Sugannasil, 2023; Narayan et al., 2022).
With the growing access to open data sources and data markets, there has been a higher demand for matching entities with external databases. Often, external databases include entities with varying levels of detail and granularity, making it rare to find an exact match. Consequently, analysts usually settle for matches with related, but not identical, entities to accomplish tasks at hand. Besides fuzzy terms, a significant challenge in such an entity matching process lies in the ambiguity in defining what ”match” means. We illustrate such ambiguity through a real-world problem:
(a) Using the ada-002 Embedding and its Euclidean distance to rank the similarity of entities. However, many entities from table $R_{2}$ are related to those in $R_{1}$ in various ways that overwhelm analysts.
(b) Using GPT-4 for Entity Matching raises ambiguity in what ”match” means. If ”match” means ”exactly the same”, it is too strict to be practically useful as none of the entities are matched.
(c) Discover the relations between source and target entities to disambiguate entity matching. Even if there are no entity with exactly the same, analysts settle for entities related in some predefined ways.
Figure 1. Entity Matching for ESG emission factor.
Example 1.1.
ESG FLO is a startup that provides auditable environmental, social, and governance (ESG) data. Their analysts use a customer-provided table $R_{1}$ of item invoices to estimate carbon emissions. Additionally, analysts use an external table, $R_{2}$, which offers ESG emissions factor estimates for various items. The challenge arises because the items listed in $R_{1}$ and $R_{2}$ are related but rarely identical. Consider the ”Charger for Consumer Electronics” item in $R_{1}$:
  • Embedding: When using the ada-002 embedding to find semantically similar items, analysts often face an overwhelming variety of items that appear related (see Figure 1(b)) but hard to match. They also find the euclidean distance hard to interpret.
  • GPT: When using the GPT-4 to decide entity matching, GPT-4 finds the term ”match” ambiguous. If ”match” means ”exactly the same”, this definition is overly strict to be practically useful, as external databases rarely have ”exactly the same” entities.
As a result, considerable manual effort is required for analysts to manually understand their relationships and determine their relevance to the task. For example, they find ”Charger for Smartphone” to be a more specific entity and ”Power Adapter” could be more general. After careful consideration, they prefer using ”Power Adapter” for emission estimation as it’s related but makes fewer assumptions, even though it’s not an exact match and not the closest in the embedding space.
The issue of ambiguous entity is widespread (Huang et al., 2023; Veltri et al., 2023). For instance, in the realm of ontologies such as OWL (Web Ontology Language) (McGuinness et al., 2004) and Schema.org (Guha et al., 2016), as well as in standards like the OHDSI Vocabularies (Bathelt, 2021), entities are often noisy and exhibit a lack of uniformity in terms of the levels of detail and granularity. To effectively match entities, it’s necessary not only to identify semantically similar entities but also to understand their relationships to the input entities, to identify the most appropriate one for the given task.
This paper proposes a novel approach to disambiguate entity matching by identifying a set of ”relations” that analysts pre-define as important for their task. Our primary observation is that the entity matching process in practice is typically iterative, rather than a straightforward one-time process. Analysts often have a predefined list of relations relevant to their task in mind. They start by seeking entities that ”exactly the same.” If such matches are not found, they may consider entities that are conceptually ”similar but differ in details” for their estimation. Throughout the process, ”relation” is the crucial factor for decision making. While the notion of relation has long been fundamental in understanding connections between entities and has been used in knowledge extraction, we are the first in applying it to disambiguate entity matching.
Example 1.2.
Continuing with the previous examples, when no exact match for the input entity, ”Charger for Consumer Electronics”, is found, analysts explore the different relations with target entities. ”Power Adapter” represents a broad category without extra details, and ”Charger for Smartphone” suggests a more specific category that assumes the electronic item is a smartphone—an assumption analysts prefer not to make. Meanwhile, entities like ”Cables and Connectors” are components of the input entity, but are considered too distant in relation. Therefore, analysts choose ”Power Adapter” as the match for estimating carbon emissions.
2. Approach Overview
2.1. Problem Definition
We follow the standard data model, where tables are denoted as $R$, comprising a set of tuples $t$, and a list of attributes $A$. Given two tables $R_{1}$ and $R_{2}$, each entry is defined by attribute-value pairs. Traditional Entity Matching (EM) looks for a function $f$:
$f:R_{1}\times R_{2}\rightarrow\{0,1\}$
This function $f$ identifies if a pair $(t_{1},t_{2})$ ”matches,” indicating they refer to exactly the same entity, by giving $f(t_{1},t_{2})=1$; otherwise, it yields $f(t_{1},t_{2})=0$.
However, in practice, finding exactly the same matches in external databases is rare, so users also seek entities that are related in different ways. With a domain of relation concepts $REL$, relation-based EM seeks a matching function $f$:
$f:REL\times R_{1}\times R_{2}\rightarrow\{0,1\}$
For every trio $(rel,t_{1},t_{2})$ where $rel\in REL$ denotes a type of relationship (such as ”Is a”, ”Contains”), $f(rel,t_{1},t_{2})=1$ if $t_{1}$ and $t_{2}$ are related by $rel$, and $f(rel,t_{1},t_{2})=0$ otherwise. The relation-based EM generalizes traditional EM, as the ”exactly the same” can be considered as just one of the many possible relations.
Remark: (1) Relations are not mutually exclusive. One entity can be related to another in multiple relations. (2) Some relations, like ”contains,” are one-to-many. For $\{t|(rel,t_{1},t)=1\land t\in R_{2}\}$, while each element is contained by $t_{1}$, the set does not indicate a complete set of components. (3) When $t_{1}$ is related to many entities for a given $rel$, how choose the best match among them depends on the relation and may require a manual process. We consider it as a postprocessing step of relation-based EM.
Figure 2. System Design.
2.2. System Design and Usage Walkthrough
This section provides a walkthrough of the system design and its usage, which includes both offline and online phases.
2.2.1. Offline
Offline, analysts decides the relations of entities that are of interest for the task, and the preprocess the external table to accelerate online entity matching.
Relation Specification. Analysts define a set of relations that are pertinent to the task during the offline brainstorming sessions. To identify these relations, analysts typically first perform entity matching manually. Then, they analyze the patterns and commonal relations useful for the task. The specificity of relation specifications is crucial, often enhanced by using examples in a few-shot learning context. This process is iterative, involving brainstorming and verification, and the relations may improve over time.
Example 2.1.
For ESG Flo, after thorough discussions regarding the requirements, the analysts agreed upon the following relations:
  • Exactly the same: For the same entities, but with synonyms or abbreviations. E.g., ”small automobile” is a synonym for ”small car”.
  • General without additional details: For entities that are a general superclass of the input entity without additional details. E.g., ”small vehicle” and ”car” as general categories for ”small car”.
  • Similar with Additional Details: For entities that include additional assumptions or features. E.g., ”electric car” adds an assumption of electricity that ”small car” does not imply.
  • Similar with wrong Details: For similar entities but with details that contradict the input entity. E.g., ”big car” contradicts the detail ”small” in ”small car” but they are both ”car”.
  • Component: For parts or constituents of the given entity. E.g., ”engine” as a component of a ”small car”.
Embedding. In the spirit of blocking, embeddings are utilized to identify generally related entities. Given an external database for matching entities, we compute the embedding of the concatencation of attributes of its rows to expedite the discovery of relations during the online phase. By default, the ard-002 embedding model is employed, with the generated embeddings being stored in a Faiss index for efficient retrieval.
2.2.2. Online Phase
In the online phase, when provided with a user table, we perform entity matching, discover relations, and carry out post-processing according to the analysts’ specific tasks.
Retrieval Augmented Generation. We generate prompts that for each tuple $t\in R_{1}$ and each specified relation $rel$. Given $t$, we retrieve a set of $K$ entities from $R_{2}$ that are near $t$ in the embedding space. The prompt inquires, for each of the $K$ entities, whether it’s related to $t$ by $rel$. To enhance accuracy and assist in interpretation, we employ a standard chain-of-thought process. These prompts are then processed by Large Language Model (GPT-4 by default). The following template is used:
     Task: Decide input & output entity relation.  Data: The input entity: {input_entity_row}  The output entities: {output_entity_rows}  Relation: {relation_description_with_example}  Steps:  1. Repeat input entity and relation.  2. Go through each output entity.  Reason if it has the relation to input entity.
The variable $K$ serves as a hyperparameter, balancing a trade-off: retrieving more entities improves recall but also increases computational demand and time. By default, we set $K=10$ and continue to retrieve the next $K$ entities if $30\%$ of the entities from the last batch are confirmed to have the specified relation.
Presentation and Post Processing. After identifying the relations, we present the results to the analyst for post-processing. This step is iterative, beginning with matches that are ”exactly the same”. If a suitable match is not found, the process continues down a predefined list of relations for the next best estimation.
Example 2.2.
For ESG Flo, the system generates a report. Each entity is first checked for an ”exactly the same” match. If found, this match is used. If not, analysts try to use ”General without additional details” as these are less assumptive. Among these, the most specific entity is selected. Should entities with ”Additional Details” be the next best match, these are then cross-referenced with customer data to ascertain the most suitable match. In cases where ”Similar with wrong Details” relations are found, efforts are made to identify the closest match. Lastly, if the only match is ”Component,” the analysts aim to compile the most significant components and aggregates their data to estimate carbon emissions.